OpenStudy (anonymous):

Elizabeth has worked for 5 months and earned 6 vacation days. If she continues to earn vacation days at this rate, how many more months does she need to work to reach a total of 10 vacation days?

OpenStudy (anonymous):

multiply 5 by 6/10 = 30/10= 3 more months